company XY uses perpetual inventory method. January's entries can be seen below:
Jan 1 Opening 300 for @ 2.30
Jan 12 Shopping 400 for @ 2.10
Jan 18 Sales 500 for @ 3.80
Jan 21 Shopping 300 for @ 2.40
Jan 25 Purchase 100 for @ 2.20
Jan 31 Sales 450 for @ 3.80

Assume that company XY uses the FIFO method, what is the value of the inventory after the sale on January 31?

Group of answer choices

A. 340

B. 1.110

C. 1.020

D. 1.005
Answer #1

Total available units = 300+400+300+100 = 1100

Units sold = 500+450 = 950

Units in inventory = 1100 - 950 = 150

Ad company uses Fifo, these units belong to last purchases

= (100*2.20) + (50*2.40)

= 340

Option A